1. The value of KC for the thermal decomposition of hydrogen sulfide, shown below, is 2.2 × 10-4 at 1400 K.
2H2S(g) ⇌ 2H2(g)+ S2(g)
A sample of gas in which [H2S] = 4.45 M is heated to 1400 K in a sealed vessel. After chemical equilibrium has been achieved, what is the value of [H2S]? Assume no H2 or S2 was present in the original sample.answer in M
1
Expert's answer
2020-04-06T11:48:30-0400

"Kc = {\frac {[H_2]^2[S_2]} {[H_2S]^2}}"

Initial concentration of H2S is 4.45 M. If 2x mol of H2S were involved in reaction, 2x mol of H2 and x mol of S2 were produced. Then in equilibrium state we have:

x mol S2

2x mol H2

4.45 - 2x mol H2S

"Kc = {\frac {[H_2]^2[S_2]} {[H_2S]^2}}={\frac {(2x)^2(x)} {(4.45-2x)^2}}=2.2*10^{-4}"

"4x^3 = 2.2*10^{-4}*(4.45-2x)^2"

After solving the equation (i did it on computer) we got x = 0.0998, then amount of H2S involved in reaction is 2x = 2*0.0998 = 0.1996, in equilibrium [H2S] = 4.45 - 0.1996 = 4.2504 M